# Grounding AI Explanations in Experience: A Reflective Cognitive Architecture for Clinical Decision Support

### v1

@@ Abstract
Effective disease prediction in modern healthcare demands the twin goals of high accuracy and transparent, clinically meaningful explanations. Existing machine learning and large language model (LLM) based approaches often struggle to balance these goals. Many models yield accurate but unclear statistical outputs, while others generate fluent but statistically unsupported narratives, often undermining both the validity of the explanation and the predictive accuracy itself. This shortcoming comes from a shallow interaction with the data, preventing the development of a deep, detailed understanding similar to a human expert’s. We argue that high accuracy and high-quality explanations are not separate objectives but are mutually reinforcing outcomes of a model that develops a deep, direct understanding of the data. To achieve this, we propose the Reflective Cognitive Architecture (RCA), a novel framework that coordinates multiple LLMs to learn from direct experience. RCA features an iterative rule refinement mechanism that improves its logic from prediction errors and a distribution-aware rules check mechanism that bases its reasoning in the dataset’s global statistics. By using predictive accuracy as a signal to drive deeper comprehension, RCA builds a strong internal model of the data. We evaluated RCA on one private and two public datasets against 22 baselines. The results demonstrate that RCA not only achieves state-of-the-art accuracy and robustness with a relative improvement of up to 40% over the baseline but, more importantly, leverages this deep understanding to excel in generating explanations that are clear, logical, evidence-based, and balanced, highlighting its potential for creating genuinely trustworthy clinical decision support systems. The code is available at https://github.com/ssssszj/RCA. Access to the CRT dataset requires an application and approval process. If approved, we will anonymize and open-source the data.
@@ 1 Introduction
Our key insight is that high predictive accuracy and high-quality explanations are not conflicting goals but are two results of a single, deeper process: developing a direct, experience-based understanding of the data. This is like how a human expert dives deep into data before drawing conclusions. We rethink predictive accuracy not just as an end goal, but as a crucial reward signal that drives the model to build a more robust and fundamental "experience" with the underlying patterns. By optimizing for correct predictions and improving robustness against the data noise common in medical datasets, the model is forced to achieve a deep data understanding. This deep understanding is the necessary condition for generating explanations that are insightful, reliable, and clinically useful, with a high-performance predictive model appearing as a valuable, simultaneous output.
To achieve this, we propose the Reflective Cognitive Architecture (RCA), a framework designed to enable LLMs to learn directly from data through a process of experience and reflection. RCA  includes two core mechanisms. The iterative rule refinement mechanism allows the model to learn from its mistakes, treating each incorrect prediction as an experience to be turned into abstract rules, thereby creating sound logical argumentation (LA). In addition to this, the distribution-aware rules check mechanism makes sure these rules are statistically based in the overall data distribution, promoting evidence-based reasoning (EBM) and reducing cognitive biases (CB).
@@ 2.1 Explainability in Disease Prediction
The quest for explainability in disease prediction is not new, but its definition and relationship with accuracy have changed over time (Sun et al., 2024). Early approaches favored models that were naturally understandable. For instance, methods like linear regression (Hoerl & Kennard, 2000; Tibshirani, 1996) and decision trees (Breiman, 2001; Prokhorenkova et al., 2018) were valued for their transparency and provided reasonable performance. However, their explanations were limited to statistical outputs, not narrative descriptions, which require significant expert analysis and place a high cognitive load on physicians.
Subsequently, more complex deep learning models emerged, achieving very high accuracy using Transformer (Hollmann et al., 2022) or FFN (Gorishniy et al., 2023). But most of them were often "black boxes," making their reasoning difficult to understand. This shifted the focus to post-hoc explanation techniques. However, these can be unfaithful to the model’s actual reasoning process. The emergence of LLMs offered a path toward generating natural language explanations. Initial efforts used LLMs to interpret statistical outputs or directly analyze tabular data. While these methods can produce fluent text, they often suffer from a disconnect from the data’s statistics. This lack of grounding is a critical weakness, as it frequently leads to a dual failure: the explanations are not only invalid and untrustworthy, but the shallow reasoning process also harms the accuracy of the prediction itself. Our work addresses this by creating a direct link between the learning process for accuracy and the generation of explanations, ensuring they are two sides of the same coin.
@@ 3.1 RCA Overview
Figure 1: RCA pipeline. RCA  uses reflective cycles and additional checks to directly analyze data, building a deep understanding that enhances both prediction accuracy and the generation of detailed, grounded explanatory statements.
The RCA pipeline, shown in Figure 1, is engineered to systematically build and refine a deep data understanding, which is the necessary precondition for achieving high accuracy and generating high-quality explanations. The process unfolds as follows:
@@
Distribution Extraction. To ensure the model’s reasoning is grounded in Evidence-based Medicine (EBM) and to mitigate Cognitive Biasing (CB), we provide it with a global statistical context. Along with data transformation in Figure 1, we extract the data distribution $\mathcal{D}_{train}$ from the training set, which summarizes the statistical properties of the entire patient cohort (e.g., means, quantiles, frequencies). This global context prevents the model from over-interpreting individual data points.
@@
Training and Prediction Phases. The training phase is a dynamic process where the rule base is iteratively refined by two other LLMs, $M_{ref}$ and $M_{chk}$, based on prediction outcomes. In the testing phase, marked with orange lines in Figure 1, the final, optimized rule base $R^{f}$ is used to generate predictions and explanations $\hat{\mathcal{Y}}^{test}$ for unseen data. The resulting explanation is thus a product of a deep, iterative learning process that was driven by the pursuit of accuracy.
@@ 3.2 Iterative Rules Optimization: Building Logical Argumentation from Experience
where $T$ is the error batch capacity. This aggregation of incorrect cases highlights deficiencies in the current rule base.
@@ 3.3 Distribution-aware Rules Check: Grounding Logic in Evidence
The iterative process ($M_{pred}$, $M_{ref}$) builds the core logical structure (LA) from the experience of pursuing accuracy, while the check mechanism ($M_{chk}$) ensures this structure is statistically robust and evidence-based (EBM, CB). This dual-process architecture ensures that RCA develops a deep, reliable understanding of the data, which is the essential foundation for achieving both high accuracy and generating trustworthy explanatory statements.
@@ 4.1 Setup
More details of datasets are provided in Appendix A.4.
Baselines. We compare RCA against 22 baselines representing diverse approaches.
  • Traditional ML Models: We include Lasso regression (Tibshirani, 1996) and Catboost (Prokhorenkova et al., 2018). These models are standard for tabular data but produce statistical artifacts (e.g., coefficients, feature importance) rather than narrative explanations, representing a baseline for expert-driven interpretation. A ‘Qwen3-235B‘ is used to generate an explanation for these models; we provide a detailed introduction in Appendix A.5.
  • LLM-based Agents: We include two agent paradigms that reflect the state-of-the-art in LLM-driven analysis, comprising 9 and 1 methods, respectively. The LLM+Tools approach equips models with predefined functions, while the LLM+Code approach utilizes a code interpreter (OpenAI, 2023). These baselines test the hypothesis that tool use abstracts away the fine-grained data interaction necessary for deep understanding. For details about the functions and code, please refer to the Appendix A.6.
@@
  • Metrics for Explanation Quality: To directly measure the quality of the generated descriptive statements, we developed four criteria grounded in cognitive science and medical practice: Cognitive Load (CL), Logical Argumentation (LA), Evidence-based Medicine (EBM), and Cognitive Biasing (CB). These criteria have been recognized by 3 doctors as clinically valuable, as they align with real-world medical assessment needs for evaluating the clarity, rationality, and reliability of explanatory content. Then doctors are invited to score each explanation on a scale of 1 to 10 for each criterion based on a detailed rubric. The rubric of these metrics are provided in Appendix A.7, while specific examples can be found in Appendix A.8.
Implementation Details. For our method, we implement RCA using both ‘Qwen2.5-72B-Instruct’ and ‘GPT-4.1-2025-04-14’ as the base LLMs(abbreviated as RCA+Qwen2.5 and RCA+GPT-4.1 to demonstrate its architectural benefits. The error batch capacity T is set to 25. The model is trained for 15 epochs on the CRT and Diabetes datasets, and 25 epochs on the Heart Disease dataset. All prompt templates used in RCAcan be found in Appendix A.12
@@ 4.2 Main Results: Deep Understanding Yields Synergistic Gains
In this section, we conduct qualitative analysis of main experiment on the CRT dataset, while the quantitative results of all datasets can be found in the Appendix A.9. Specifically, the qualitative analysis includes Lasso, Catboost, ‘DeepSeek-Chat-V3.1‘(abbreviated as ‘DeepSeek-V3.1‘), ‘o4-mini-2025-04-16’(abbreviated as ‘o4-mini’), ‘Qwen3-235B-A22B-Instruct-2507’ (abbreviated as ‘Qwen3-235B’) and ‘o3-mini-2025-01-31+Code’ (abbreviated as ‘o3-mini+Code’) as baselines .
Our central thesis posits that predictive accuracy and explanation quality are not in opposition but are synergistic outcomes of a model’s deep, first-hand understanding of data. We test this by plotting model performance against explanation quality. As shown by the hollow points in Figure 2, which represent the main experimental results, the RCA-based approaches consistently occupy the top-right quadrant, achieving the highest predictive performance (Accuracy and MCC) while also delivering high-quality explanations (as indicated by the Cognitive Load score, with detailed metrics in Appendix A.9).
@@
Ultimately, the results show that by prioritizing the development of a deep data understanding, RCA  organically achieves state-of-the-art results in both prediction and explanation, demonstrating their synergistic relationship. Full results for all baselines across three datasets please refer to Appendix A.9.
@@ 4.3 Robustness: Deep Understanding Confers Resilience
Similarly, we conduct qualitative analysis of robust experiment on the CRT dataset, while the quantitative results can be found in the Appendix A.10.
@@
The results, visualized by the solid points and connecting dashed lines in Figure 2, demonstrate RCA’s superior robustness. The performance of RCA  (both ‘Qwen2.5-72B’ and ‘GPT-4.1’ versions) shows minimal degradation across all noise conditions, as indicated by the short dashed lines connecting the hollow (original) and solid (noisy) points. This stability is a direct outcome of its design. The distribution-aware rules check grounds the model’s logic in global statistics, preventing it from being misled by local anomalies or outliers. The iterative rules optimization builds a generalized understanding from cumulative experience, making the model less dependent on any single data point or feature.
@@ 4.4 Ablation Study
To validate that RCA’s performance stems directly from its proposed cognitive mechanisms, we conducted an ablation study by systematically removing its core components. As shown in Table 1, the removal of any key module leads to a significant performance collapse, confirming that each part is essential to the process of building a deep understanding. More explanation results are provided in Appendix A.11.
@@ 5 Conclusion
In this paper, we challenged the conventional trade-off between predictive accuracy and explainability in clinical AI. We argued that these are not competing goals but synergistic outcomes of a model that develops a deep, first-hand understanding of the data. To achieve this, we introduced RCA, a novel framework that learns from experience through iterative rules optimization and grounds its reasoning in global statistics via a distribution-aware check. Our experiments demonstrate that by forcing the model to achieve this deeper comprehension, RCA  not only attains state-of-the-art accuracy and robustness but also excels in generating clear, logical, and evidence-based explanatory statements.
@@ Appendix A Appendix / supplemental material
A.1 Declaration on the Use of LLMs
We declare that the use of LLMs during the preparation of this manuscript was strictly limited to language-related assistance, such as sentence refinement and grammatical correction. All substantive content was independently authored by the authors and rigorously reviewed and verified following any LLM-assisted modifications. During the experiments, all usage of LLMs was solely for academic research purposes, with no inappropriate applications. Detailed experimental settings are provided in the Experiments section of this paper. No other reliance on LLMs is involved in this work.
A.2 Ethics Statement
All authors of this paper strictly adhere to the ICLR Code of Ethics throughout the entire research and manuscript preparation process. Prior to submitting this work, every author has carefully read and fully understood the content of the Code of Ethics, and confirms compliance with all its provisions.
This research strictly follows the ethical guidelines for conference participation outlined in the ICLR Code of Ethics, covering the integrity of paper submission, adherence to ethical standards in potential peer review , and constructive, respectful communication in any subsequent academic discussions related to this work. We further confirm that no content in this paper violates the ethical principles specified in the ICLR Code of Ethics, and that all research procedures are conducted with honesty, transparency, and respect for academic ethics.
A.3 Reproducibility Statement
The work in this paper is well reproducible. The code used in the study is available at https://anonymous.4open.science/r/anonym107. Among the three datasets employed in the experiments, the Diabetes Dataset(https://www.kaggle.com/datasets/nanditapore/healthcare-diabetes/data) and Heart Disease Dataset(https://www.kaggle.com/datasets/oktayrdeki/heart-disease) are public and can be accessed via the corresponding links; only the CRT Dataset is private, and its access requires submission of an application and subsequent approval. Data processing can refer to Section 3.1.
A.4 Dataset Details

### v2

# Rethinking Explainable Disease Prediction: Synergizing Accuracy and Reliability via Reflective Cognitive Architecture

@@ Abstract
In clinical decision-making, predictive models face a persistent trade-off: accurate models are often opaque "black boxes," while interpretable methods frequently lack predictive precision or statistical grounding. In this paper, we challenge this dichotomy, positing that high predictive accuracy and high-quality descriptive explanations are not competing goals but synergistic outcomes of a deep, first-hand understanding of data. We propose the Reflective Cognitive Architecture (RCA), a novel framework designed to enable Large Language Models (LLMs) to learn directly from tabular data through experience and reflection. RCA integrates two core mechanisms: an iterative rules optimization process that refines logical argumentation by learning from prediction errors, and a distribution-aware rules check that grounds this logic in global statistical evidence to ensure robustness. We evaluated RCA against over 20 baselines—ranging from traditional machine learning to advanced reasoning LLMs and agents—across diverse medical datasets, including a proprietary real-world Catheter-Related Thrombosis (CRT) cohort. Crucially, to demonstrate real-world scalability, we extended our evaluation to two large-scale datasets. The results confirm that RCA achieves state-of-the-art predictive performance and superior robustness to data noise while simultaneously generating clear, logical, and evidence-based explanatory statements, maintaining its efficacy even at scale. The code is available at https://github.com/ssssszj/RCA.
Machine Learning, ICML
@@ 1 Introduction
Our key insight is that high predictive accuracy and high-quality explanations are not conflicting goals but are two results of a single, deeper process: developing a direct, experience-based understanding of the data. This is like how a human expert dives deep into data before drawing conclusions. We rethink predictive accuracy not just as an end goal, but as a crucial reward signal that drives the model to build a more robust and fundamental "experience" with the underlying patterns. By optimizing for correct predictions and improving robustness against the data noise common in medical datasets, the model is forced to achieve a deep data understanding. This deep understanding is the necessary condition for generating explanations that are insightful, reliable, and clinically useful, with a high-performance predictive model appearing as a valuable, simultaneous output. The core objective is to predict and explain, but robustness to data noise should also be considered for real-world clinical applications.
To achieve this, we propose the Reflective Cognitive Architecture (RCA), a framework designed to enable LLMs to learn directly from data through a process of experience and reflection. RCA  includes two core mechanisms. The iterative rule refinement mechanism allows the model to learn from its mistakes, treating each incorrect prediction as an experience to be turned into abstract rules, thereby creating sound logical argumentation (LA). In addition, a distribution-aware rules check mechanism grounds these rules in the statistical reality of the training data. This mechanism uses a summary of the data distribution as a contextual "sanity check.", promoting evidence-based medicine (EBM) and reducing cognitive biases (CB).
@@ 2.1 Explainability in Disease Prediction
The quest for explainability in disease prediction is not new, but its definition and relationship with accuracy have changed over time (Sun et al., 2024). Early approaches favored models that were naturally understandable. For instance, methods like linear regression (Hoerl and Kennard, 2000; Tibshirani, 1996) and decision trees (Breiman, 2001; Prokhorenkova et al., 2018) were valued for their transparency and provided reasonable performance. However, their explanations were limited to statistical outputs, not narrative descriptions, which require significant expert analysis and place a high cognitive load on physicians.
Subsequently, more complex deep learning models emerged, achieving very high accuracy using Transformer (Hollmann et al., 2022) or FFN (Gorishniy et al., 2023). But most of them were often "black boxes," making their reasoning difficult to understand. This shifted the focus to post-hoc explanation techniques, which can be unfaithful to the model’s actual reasoning process. Concurrently, Neural-Symbolic (NeSy) Networks like LTN(Badreddine et al., 2022) and LNN(Riegel et al., 2020) were explored to integrate logic into neural networks. While inherently transparent, they often struggle to generate narrative explanations or automatically discover complex rules from noisy data.
The emergence of LLMs, particularly those specialized for the medical domain (Medical LLMs) (Dou et al., 2025), offered a path toward generating natural language explanations. Initial efforts used LLMs to interpret statistical outputs or directly analyze tabular data. While these methods can produce fluent text, they often suffer from a disconnect from the data’s statistics. This lack of grounding is a critical weakness, as it frequently leads to a dual failure: the explanations are not only invalid and untrustworthy, but the shallow reasoning process also harms the accuracy of the prediction itself. Furthermore, an explanation is still insufficient even if it is fluent and statistically grounded; in high-stakes settings like healthcare, it must be demonstrably usable by and useful to human experts (Slack et al., 2023).Our work addresses the question by creating a direct link between the learning process for accuracy and the generation of explanations, ensuring they are two sides of the same coin.
@@ 3.1 RCA Overview
Figure 1: RCA pipeline. $S^{error}$ is a collection of misclassified samples (errors) from $M_{pred}$. $R^{k}$ is the rule base for the current iteration $k$, and $\hat{R^{k}}$ denotes the checked and refined rule base from iteration $k$. RCA  uses reflective cycles and additional checks to directly analyze data, building a deep understanding that enhances both prediction accuracy and the generation of detailed, grounded explanatory statements.
As illustrated in Figure 1, the RCA  pipeline systematically builds and refines data understanding. The process unfolds as follows:
@@
Distribution Extraction. To ensure the model’s reasoning is grounded in Evidence-based Medicine (EBM) and to mitigate Cognitive Biasing (CB), we provide it with a global statistical context. Along with data transformation in Figure 1, we extract the data distribution $\mathcal{D}_{train}$ from the training set, which summarizes the statistical properties of the entire patient cohort (e.g., means, quantiles, frequencies). This global context prevents the model from over-interpreting individual data points. A concrete example of such a distribution summary is provided in  A.12.
@@
Dynamic Maintenance of the Rule Base. The rule base $R$ is maintained through a dual-process feedback loop. The iterative rules optimization mechanism functions as a "reflective optimizer," refining logic based on instance-level prediction errors, while the distribution-aware rules check mechanism acts as a "statistical validator," auditing these updates against global evidence ($\mathcal{D}_{train}$) . This cyclic interaction ensures $R$ evolves dynamically at each epoch, balancing the precision required to correct specific failures with the robustness needed to prevent overfitting.
Training and Prediction Phases. The training phase is a dynamic process where the rule base is iteratively refined by two LLMs, $M_{ref}$ and $M_{chk}$, based on prediction outcomes. In the testing phase, marked with orange lines in Figure 1, the final, optimized rule base $R^{f}$ is used to generate predictions and explanations $\hat{\mathcal{Y}}^{test}$ for unseen data. The resulting explanation is thus a product of a deep, iterative learning process that was driven by the pursuit of accuracy.
@@ 3.2 Iterative Rules Optimization: Building Logical Argumentation from Experience
where $T$ is the error batch capacity, and $\mathrm{conc}$ is concatenation function that stitches samples together. This aggregation of incorrect cases highlights deficiencies in the current rule base.
@@ 3.3 Distribution-aware Rules Check: Grounding Logic in Evidence
The iterative process ($M_{pred}$, $M_{ref}$) builds the core logical structure (LA) from the experience of pursuing accuracy, while the check mechanism ($M_{chk}$) ensures this structure is statistically robust and evidence-based (EBM, CB). This dual-process architecture ensures that RCA develops a deep, reliable understanding of the data, which is the essential foundation for achieving both high accuracy and generating trustworthy explanatory statements. To illustrate this synergistic process, a detailed walkthrough of the entire reflective cycle is presented in  A.1.
@@ 4.1 Setup
More details of datasets are provided in  A.2.
Furthermore, to thoroughly test the real-world scalability of our approach, we included two additional, large-scale datasets in our evaluation, CRT_ex and Cardiovascular Disease (Shihab, 2025). The experimental setup for these datasets followed the same protocol as our main experiments. To maintain focus on our core contributions, the detailed results and analysis for this scalability study are presented in  A.10.
Baselines. We compare RCA against 25 baselines representing diverse approaches, including Traditional ML Models, Neural-symbolic Networks, LLM-based Methods, Reasoning LLMs, Medical LLM and LLM-based Agents. All LLM baselines were provided with the same global data distribution ($\mathcal{D}_{train}$) and illustrative in-context examples from the training set that RCA  utilized. For each baseline, we provide a detailed introduction in  A.3
@@
  • Metrics for Explanation Quality: To directly measure the quality of the generated descriptive statements, we developed four criteria grounded in cognitive science and medical practice: Cognitive Load (CL), Logical Argumentation (LA), Evidence-based Medicine (EBM), and Cognitive Biasing (CB). Details of these metrics are provided in  A.4, while specific examples can be found in  A.7.
  • Metrics for Explanation Usability:To further validate usability, we conducted a formal human-centric study to assess practical usability. We asked 3 expert clinicians to review 50 sample explanations from each of the three datasets.In this study, participants compared the explanation from RCA against the explanation from strongest baseline ‘Qwen3-235B’. For each pair, the clinicians were asked to identify the superior explanation across our four core criteria. This forced-choice design allows us to directly and formally quantify the expert preference for RCA’s explanations. Analysis is provided in  A.8
Implementation Details. For our method, we implement RCA using both ‘Qwen2.5-72B-Instruct’ and ‘GPT-4.1-2025-04-14’ as the base LLMs(abbreviated as RCA+Qwen2.5 and RCA+GPT-4.1 to demonstrate its architectural benefits. The error batch capacity T is set to 25. The model is trained for 15 epochs on the CRT and Diabetes datasets, and 25 epochs on the Heart Disease dataset. All prompt templates used in RCAcan be found in  A.13
@@ 4.2 Main Results: Deep Understanding Yields Synergistic Gains
In this section, we conduct qualitative analysis of main experiment on the CRT dataset, while the quantitative results of all datasets can be found in the  A.8. Specifically, the qualitative analysis includes Lasso, Catboost, ‘DeepSeek-Chat-V3.1‘(abbreviated as ‘DeepSeek-V3.1‘), ‘o4-mini-2025-04-16’(abbreviated as ‘o4-mini’), ‘Qwen3-235B-A22B-Instruct-2507’ (abbreviated as ‘Qwen3-235B’) and ‘o3-mini-2025-01-31+Code’ (abbreviated as ‘o3-mini+Code’) as baselines .
Our central thesis posits that predictive accuracy and explanation quality are not in opposition but are synergistic outcomes of a model’s deep, first-hand understanding of data. We test this by plotting model performance against explanation quality. As shown by the hollow points in Figure 2, which represent the main experimental results, the RCA-based approaches consistently occupy the top-right quadrant, achieving the highest predictive performance (Accuracy and MCC) while also delivering high-quality explanations (as indicated by the Cognitive Load score, with detailed metrics in  A.8).
@@
Ultimately, the results show that by prioritizing the development of a deep data understanding, RCA  organically achieves state-of-the-art results in both prediction and explanation, demonstrating their synergistic relationship. Full results for all baselines across three datasets please refer to  A.8.
@@ 4.3 Robustness: Deep Understanding Confers Resilience
Similarly, we conduct qualitative analysis of robust experiment on the CRT dataset, while the quantitative results can be found in the  A.9.
@@
The results, visualized by the solid points and connecting dashed lines in Figure 2, demonstrate RCA’s superior robustness. The performance of RCA  (both ‘Qwen2.5-72B’ and ‘GPT-4.1’ versions) shows minimal degradation across all noise conditions, as indicated by the short dashed lines connecting the hollow (original) and solid (noisy) points. This stability is a direct outcome of its design. The distribution-aware rules check grounds the model’s logic in global statistics, preventing it from being misled by local anomalies or outliers. This mechanism is particularly effective at identifying and flagging extreme outliers that might otherwise corrupt predictions, as detailed in the edge case analysis in  A.5. The iterative rules optimization builds a generalized understanding from cumulative experience, making the model less dependent on any single data point or feature.
@@ 4.4 Ablation Study
To validate that RCA’s performance stems directly from its proposed cognitive mechanisms, we conducted an ablation study by systematically removing its core components. As shown in Table 1, the removal of any key module leads to a significant performance collapse, confirming that each part is essential to the process of building a deep understanding. More explanation results are provided in  A.11.
@@ 5 Conclusion
In this paper, we argued that the conventional trade-off between predictive accuracy and explainability are not competing goals, but synergistic outcomes of a model that develops a deep understanding of the data. To achieve this, we introduced RCA, a novel framework that learns from experience and anchors its reasoning in global statistics. Our experiments demonstrate that by forcing the model to achieve this deeper comprehension, RCA  not only attains state-of-the-art accuracy but also excels in generating clear, logical, and evidence-based explanatory statements.
@@ Appendix A Appendix / supplemental material
A.1 RCA REFLECTIVE PROCESS: A WALKTHROUGH
To provide a concrete illustration of the RCA framework, this section details the interaction between its two core mechanisms: Iterative Rules Optimization ($M_{ref}$) and Distribution-aware Rules Check ($M_{chk}$). We trace how the system processes a specific patient record during the training phase to demonstrate the synergy between instance-level learning and global-level statistical grounding.
A.1.1 Step 1: Initial State (Context)
Consider a scenario where the model is in the middle of the training process (Epoch $k$). It has already acquired some feature-coupled logic, but the rule base ($R^{k-1}$) lacks precision for borderline cases.
  • Current Rule Base ($R^{k-1}$): Contains broad, clinically plausible rules, for example:
    • Rule A: "Patients undergoing Chemotherapy with a Central Catheter (CVC/PICC) are considered high risk if D-dimer >0.50>0.50 mg/L, reflecting the prothrombotic effect of systemic treatment."
    • Rule B: "High inflammatory status, indicated by GLR >3.0>3.0, coupled with Age >60>60, predicts thrombosis even with normal D-dimer levels."
  • Global Data Distribution ($\mathcal{D}_{train}$): The model possesses the statistical ground truth of the dataset. For instance, it identifies that for no-CRT patients in this specific cancer cohort, the 75th percentile of D-dimer is actually 1.12 mg/L (notably higher than the standard 0.50 cutoff in Rule A), and the Granulocyte-to-lymphocyte ratio (GLR) typically ranges from 0.86 to 2.63.
A.1.2 Step 2: The Prediction Error (Short-term Experience)
The prediction LLM ($M_{pred}$) encounters a specific instance, Patient 4, with the following profile:
  • Features: GLR = 0.94 (Low), D-dimer = 0.58 (Marginally elevated), Chemotherapy = Yes, Catheter = PICC.
  • True Label: No CRT.
The Error: Applying Rule A, the model identifies the risk factors (Chemotherapy + PICC) and notes the D-dimer (0.58) exceeds the 0.50 threshold. Consequently, it incorrectly predicts "CRT". This failure constitutes the "short-term experience" ($S^{error}$) passed to the reflection module.
A.1.3 Step 3: Iterative Rules Optimization (Bottom-Up Refinement)
The Reflection LLM ($M_{ref}$) analyzes this specific error. It recognizes that Rule A was too aggressive for this specific context: for Patient 4, a D-dimer of 0.58 was non-thrombotic, likely mitigated by the low inflammatory state (GLR=0.94).
To rectify this, $M_{ref}$ proposes a new, highly specific rule derived directly from the instance:
Proposed Rule (Specific): "If the patient has a GLR between 0.9–1.0 AND a D-dimer level between 0.4 and 0.6, predict No CRT, overriding risk factors like Chemotherapy."
While this rule successfully corrects the mistake by effectively "memorizing" Patient 4, it is dangerously over-fitted to the specific values of a single instance.
A.1.4 Step 4: Distribution-aware Rules Check (Top-Down Safeguard)
At the end of the epoch, the Check LLM ($M_{chk}$) validates this proposed rule against the Global Data Distribution ($\mathcal{D}_{train}$).
  • It observes that the proposed D-dimer range (0.4–0.6) falls well within the "safe zone" for the broader non-thrombosis population (where the 75th percentile is 1.12).
  • It notes that the specific GLR range (0.9–1.0) is merely a narrow slice of the typical healthy range (0.86–2.63).
Recognizing that the specific rule is a subset of a broader statistical truth, $M_{chk}$ generalizes it into a robust, evidence-based rule that replaces the over-fitted proposal:
Final Refined Rule (General): "If the patient has a GLR between 0.86–2.63 (Normal Range) AND a D-dimer level less than 1.12, they are less likely to develop CRT, even in the presence of Chemotherapy."
A.1.5 Summary
This walkthrough highlights the core philosophy of RCA: the Refinement mechanism uses specific errors to discover nuances (e.g., "Standard D-dimer cutoffs don’t apply here"), while the Check mechanism uses global statistics to validate and generalize those discoveries (e.g., "Actually, anything under 1.12 is safe for this group"). This balance prevents overfitting to noise while ensuring the final rules are both accurate and robust.
A.2 Dataset Details